OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. (2 point)

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

parallel lines has equal gradient then y=-1(x)+c 2=-1(-4)+c c=-2 then the equation is y=-x-2